# Mechanisms and Computational Design of Multi-Modal End-Effector with Force Sensing using Gated Networks

### v3

@@ V-D Grasping and Foot Capability
Here, we verify MAGPIE’s grasping and capability as a foot. MAGPIE employs a custom direct drive BLDC with rated speed and torque of $114$ rad/sec and $0.58$ N/m and peak torque of $0.96$ N/m. With the lead screw linear mechanisms, the nominal gripper opening and closing speed of $58.3$ mm/sec and the nominal grasping force of $350$ N. The rigidity of the entire mechanism was evaluated by applying a static load of $200$ N and measuring the deflection, which was $0.3$ mm.
@@ VI Conclusion
This paper presented MAGPIE, an end-effector capable of grasping an object while functioning as both line and flat feet with contact/tactile force sensing through 3D hall effect sensors. Our computational design framework and design search enable the design of the sensing mechanisms, including flexure beams, magnet, and sensor parameters, while considering nonlinear effects and magnetic disturbances, such as interference from neighboring sensor magnets when the gripper is closed. MAGPIE mechanisms and force-sensing capabilities have been verified on the hardware using an ideal GRBF model generated from the framework and stacked GRU. MAGPIE enables the development and deployment of limbed robotic systems that can leverage MAGPIE’s multi-modality and total 8-axis contact and tactile force sensing for conducting, such as simultaneous locomotion, manipulation, and grasping tasks. Our computational design framework provides a way to further develop hall effect-based force sensing mechanisms that are more resilient and less prone to self-interference. MAGPIE showcased multi-modal robotic end-effectors, promising enhanced versatility in multi-modal robotics.

### v4

@@ V-D Grasping and Foot Capability
Here, we verify MAGPIE’s grasping and capability as a foot. MAGPIE employs a custom direct drive BLDC with rated speed and torque of $114$ rad/sec and $0.58$ N/m and peak torque of $0.96$ N/m. With the lead screw mechanisms, the nominal gripper opening and closing speed was $58.3$ mm/sec, and the nominal grasping force was $350$ N, which was measured using our Hall effect sensors. The rigidity of the entire mechanism was evaluated by applying a static load of $200$ N at the top of the gripper and measuring the deflection, which was $0.3$ mm.
@@ VI Conclusion
This paper presented MAGPIE, an end-effector capable of grasping an object while functioning as both line and flat feet with contact/tactile force sensing through 3D Hall effect sensors. Our computational framework enables the design of the sensing mechanisms, including flexure beams, magnet, and sensor parameters, while considering nonlinear effects and magnetic disturbances. MAGPIE mechanisms and force-sensing capabilities have been verified on the hardware using an ideal GRBF model generated from the framework and stacked GRU. MAGPIE enables the development and deployment of limbed robotic systems that can leverage MAGPIE’s multi-modality and total eight-axis contact and tactile force sensing for conducting, such as simultaneous locomotion, manipulation, and grasping tasks.
Our multi-axis sensing mechanism can be extended and applied to smaller or larger end effector force sensing scenarios thanks to the simplicity of the mechanism and scalability provided by the computational framework. MAGPIE computational framework provides users "off-the-shelf" ideal models for customized sensing mechanisms, while allowing users to tailor the sensor model with our GRU framework once data is collected. Hence, our work will aid the broader adoption of multi-axis Hall effect sensor-based force sensing. MAGPIE showcased multi-modal robotic end-effectors, promising enhanced versatility in multi-modal robotics.